OpenStudy (anonymous):

Identify the number of solutions to the quadratic equation 2x2 – 3x + 4 = 0 A. 0 B. 1 C. 2

OpenStudy (anonymous):

you want to go ahead and solve if, or just answer the question?

OpenStudy (anonymous):

jus answer i gotta get it done asap to pass

OpenStudy (anonymous):

\[2x^2-3x+4=0\] looks like \[ax^2+bx+c=0\] where \(a=2,b=-3,c=4\) compute \(b^2-4ac=(-3)^2-4\times 2\times 4\) and tell me what you get

OpenStudy (anonymous):

c

OpenStudy (anonymous):

it be c

OpenStudy (anonymous):

did you compute?

OpenStudy (anonymous):

i forget how to